OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 19 feet, what is its lateral area? Use π = 22 over 7 1980 ft2 836 ft2 990 ft2 1672 ft2

OpenStudy (anonymous):

Figure the radius from the area of the base, first.

OpenStudy (anonymous):

Can you?

OpenStudy (anonymous):

no can you explain how?

OpenStudy (anonymous):

What is the area of a circle, it's the same.. (this is a cylinder).

OpenStudy (anonymous):

ughh this is mkn my head hurt . im sorry ive took 4 exams today so far .

OpenStudy (anonymous):

Area = π × r2

OpenStudy (anonymous):

So 154 = 22/7 times r^2 What is r?

OpenStudy (anonymous):

to find that wud i half to multiply 22/7 * 154 then divide by 2?

OpenStudy (anonymous):

Not quite. sqrt(154 times 7/22)

OpenStudy (anonymous):

so it wud be 154 * 22/7 which is 484 then square it ? so it wud be 484^2 ?

OpenStudy (anonymous):

reread

OpenStudy (anonymous):

im sorry i dont get it . :/

OpenStudy (anonymous):

You multiplied by 22/7 instead of 7/22

OpenStudy (anonymous):

ok i got 49.

OpenStudy (anonymous):

OK, now what?

OpenStudy (anonymous):

Tell you what, let's go step by step, OK?

OpenStudy (anonymous):

yes please.

OpenStudy (anonymous):

We have pi r^2 = 154 = (22/7) r^2 OK?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So multiply both sides by 7 7 x 154 = 22 r^2

OpenStudy (anonymous):

(7 x 154)/22 = r^2 OK?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now do the multiplication, then take the square root of both sides

OpenStudy (anonymous):

The 49 is what you got before, but you need the square root.

OpenStudy (anonymous):

so next would be 1,078/22=r^2 then its 49=r^2 then 49^2 n that is 2,401

OpenStudy (anonymous):

square root

OpenStudy (anonymous):

ooohh sorry . its 7

OpenStudy (anonymous):

Yes! So r = 7. OK?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

The other part of the problem is to find the "lateral area", the area of the "skin" on the cyclinder (not counting the ends), like an open tube.

OpenStudy (anonymous):

There's a formula for that, but I want you to visualize sliding a rubber band down the length of the cylinder. Can you see that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The area the rubber band will cover is the perimeter of the circle (length of the rubber band) times how far we roll it

OpenStudy (anonymous):

So the formula is 2 pi r times the height of the cylinder.

OpenStudy (anonymous):

ok . so tht wud be 836

OpenStudy (anonymous):

2 * 22/7 * 7 * 19 = 836

OpenStudy (anonymous):

Sounds good to me.

OpenStudy (anonymous):

thank you soooo much . you helped me soo much

OpenStudy (anonymous):

yw